OpenStudy (anonymous):

PLEASE HELP ASAP... WILL REWARD AND MEDAL -what is the sales tax on $363.85 ?????

OpenStudy (mary.rojas):

well what is the tax %

OpenStudy (anonymous):

idk

OpenStudy (mary.rojas):

okay so 7%...lets see

OpenStudy (luigi0210):

uh, is there anymore information?

OpenStudy (mary.rojas):

nope only that 7%

OpenStudy (luigi0210):

Multiply them? Then add what you get

OpenStudy (mary.rojas):

but the tax is already included......in the total? smh this is giving me a headache

OpenStudy (luigi0210):

Oh, okay.

OpenStudy (mary.rojas):

wait idk she said she would be back in a few minutes

OpenStudy (anonymous):

im here

OpenStudy (mary.rojas):

ok is the tax already added in the total?

OpenStudy (anonymous):

noooo

OpenStudy (mary.rojas):

ohhh lol okayz

OpenStudy (mary.rojas):

that is what I was confused about lol!

OpenStudy (mary.rojas):

363.85*0.07=

OpenStudy (anonymous):

multiply?

OpenStudy (mary.rojas):

what is that i dont have a calculator

OpenStudy (mary.rojas):

yeashhh

OpenStudy (anonymous):

I multiplied 363.85*0.07=25.4695

OpenStudy (mary.rojas):

okay so that is your tax amount? @Luigi0210

OpenStudy (anonymous):

what tax is narrowed down?

OpenStudy (mary.rojas):

what do you mean like rounded?

OpenStudy (anonymous):

yaa

OpenStudy (mary.rojas):

so... 25.47

OpenStudy (luigi0210):

Yea, that's the tax amount, but not the total cost.

OpenStudy (mary.rojas):

yeah she just wants the tax amount

OpenStudy (anonymous):

what is the total cost now?? from subtotal-$363.85 and tax-$25.47 now what is the total???

OpenStudy (mary.rojas):

add them doll

OpenStudy (anonymous):

$363.85+25.47=389.32

OpenStudy (anonymous):

total is $389.32 ??????

OpenStudy (mary.rojas):

yeshhhh